# Magic Gems: A Polyhedral Framework for Magic Squares Connecting Combinatorics, Geometry, and Linear Algebra

### v1

@@ Abstract
We introduce Magic Gems, a geometric representation of magic squares as three-dimensional polyhedra. By mapping an $n\times n$ magic square onto a centered coordinate grid with cell values as vertical displacements, we construct a point cloud whose convex hull defines the Magic Gem. Building on prior work connecting magic squares to physical properties such as moment of inertia, this construction reveals an explicit statistical structure: we show that magic squares have vanishing covariances between position and value. We develop a covariance energy functional—the sum of squared covariances with row, column, and diagonal indicator variables—and establish rigorously for $n=3$ (via exhaustive enumeration) that its zeros correspond precisely to magic squares. Large-scale sampling for $n=4,5$ (460+ million arrangements) provides strong numerical evidence that this characterization extends to larger orders. Perturbation analysis demonstrates that magic squares are isolated local minima. The representation is invariant under dihedral symmetry $D_{4}$, yielding canonical geometric objects for equivalence classes.
@@ 1.1 Main Contributions
Our main theoretical contribution is the Energy Characterization Theorem (Section 3.13), which establishes that magic squares correspond to zeros of a covariance energy functional. We prove this result rigorously for $n=3$ via exhaustive enumeration, and provide strong computational evidence for larger orders. This result extends the physical interpretation of magic squares initiated by Loly [7], who showed that the moment of inertia tensor of a magic square (treating entries as masses) exhibits special invariance properties. Our covariance formulation makes this connection explicit and enables systematic perturbation analysis, confirming that magic squares are isolated points in the arrangement space.
@@ 1.2 Paper Organization
The remainder of this paper proceeds as follows. Section 2 reviews classical results on magic squares and establishes notation for the geometric and statistical concepts we employ. Section 3 presents the Magic Gem construction, proves the Energy Characterization Theorem, and develops the perturbation analysis framework. Section 4 describes our computational experiments, validating theoretical predictions and exploring the structure of the arrangement space for $n=3$, 44, and $5$. Section 5 interprets our findings and discusses connections to related areas. Finally, Section 6 summarizes our contributions and outlines directions for future research.
@@ 3.1 From Magic Square to Point Cloud
Section 1 illustrates the four-step construction process, progressing from the magic square through the centered grid to the final three-dimensional representation, along with all eight $D_{4}$ symmetry variants of the Lo Shu square.
Figure 1: Construction process and symmetry structure. Top row (a–d): Construction of a Magic Gem from the Lo Shu square. (a) The Lo Shu magic square—the unique (up to symmetry) $3\times 3$ magic square, with each row, column, and diagonal summing to $M(3)=15$. (b) Placement on a centered coordinate grid with origin at center. (c) Assignment of $z$-values from cell entries. (d) Centered Magic Gem point cloud with convex hull. Bottom rows (e–l): The eight $D_{4}$ symmetries of the $3\times 3$ magic square, comprising four rotations and four reflections. Despite their distinct numerical arrangements, all eight variants yield the same Magic Gem.
@@ 3.2 The Magic Gem Polyhedron
The construction figure (Section 1, bottom rows) displays all eight symmetry variants of the $3\times 3$ magic square, illustrating how these numerically distinct arrangements share a common geometric representation.
@@ 3.4.1 Forward Direction: Magic Implies Zero Covariance
3.4.2 On the Insufficiency of Row and Column Covariances Alone
It is natural to ask whether the conditions $\mathrm{Cov}(X,Z)=\mathrm{Cov}(Y,Z)=0$ alone suffice to characterize magic squares, or at least semi-magic squares (those with equal row and column sums but possibly failing the diagonal conditions). Perhaps surprisingly, the answer is no.
Remark 3.8 (Counterexamples Exist).
@@ 3.4.2 On the Insufficiency of Row and Column Covariances Alone
This observation has important implications. The covariance conditions with $X$ and $Y$ alone impose only two linear constraints on the space of possible row and column sum patterns. The remaining freedom in these patterns can be exploited to achieve zero covariance without achieving equal sums. The characterization of magic squares thus requires all four covariance conditions, not merely the row and column covariances. This motivates the energy functional approach developed in Section 3.7.
@@ 3.5 Diagonal Conditions as Covariances
This reformulation provides a unified covariance-based characterization: an arrangement is a magic square if and only if it has zero covariance with all four position indicators—$X$, $Y$, $D_{\mathrm{main}}$, and $D_{\mathrm{anti}}$.
@@ 3.6 Perturbation Analysis
This result establishes that magic squares correspond to global minima of the functional $\Phi$. Since $\Phi\geq 0$ with equality only for magic configurations (among integer arrangements), magic squares are isolated critical points in the covariance landscape. Generic perturbations—those not preserving all row and column sums—immediately increase $\Phi$, confirming the exceptional nature of magic configurations.
@@ 3.7 The Covariance Energy Landscape
Definition 3.12 (Covariance Energy Functional).
For an arrangement $S$, define the total magic energy as the sum of squared covariances with all four position indicators:
$E(S)=\mathrm{Cov}(X,Z)^{2}+\mathrm{Cov}(Y,Z)^{2}+\mathrm{Cov}(D_{\mathrm{main}},Z)^{2}+\mathrm{Cov}(D_{\mathrm{anti}},Z)^{2}.$ (12)
Equivalently, we may decompose this as
$E(S)=E_{\text{cov}}(S)+E_{\text{diag}}(S),$
where $E_{\text{cov}}(S)=\mathrm{Cov}(X,Z)^{2}+\mathrm{Cov}(Y,Z)^{2}$ captures row/column imbalance and $E_{\text{diag}}(S)=\mathrm{Cov}(D_{\mathrm{main}},Z)^{2}+\mathrm{Cov}(D_{\mathrm{anti}},Z)^{2}$ captures diagonal imbalance.
This formulation is conceptually clean: the magic energy is simply the sum of squared “correlations” between $z$-values and each of the four structural directions (rows, columns, main diagonal, anti-diagonal). An arrangement achieves zero energy precisely when it is uncorrelated with all four directions—the definition of a magic square.
Theorem 3.13 (Energy Characterization).
For $n=3$, an arrangement $S$ of $\{1,\ldots,n^{2}\}$ is a magic square if and only if $E(S)=0$. This result is proven by exhaustive enumeration.
@@
The forward direction follows from Section 3.6 and Section 3.10: every magic square has $\mathrm{Cov}(X,Z)=\mathrm{Cov}(Y,Z)=0$ (from equal row and column sums) and $\mathrm{Cov}(D_{\mathrm{main}},Z)=\mathrm{Cov}(D_{\mathrm{anti}},Z)=0$ (from equal diagonal sums), hence $E(S)=0$.
For the converse, we rely on exhaustive enumeration. Among all 9!=362,8809!=362{,}880 permutations of $\{1,\ldots,9\}$, exactly 8 arrangements achieve $E(S)=0$ (to machine precision). All 8 of these arrangements are magic squares—specifically, the Lo Shu square and its seven $D_{4}$ symmetry variants. No false positives exist: every arrangement with $E(S)=0$ is magic.
Note that the individual covariance conditions are not independent: as shown in Section 3.8, 760 arrangements satisfy $\mathrm{Cov}(X,Z)=\mathrm{Cov}(Y,Z)=0$, but only the 8 magic squares additionally satisfy the diagonal conditions. The characterization requires all four covariances to vanish simultaneously. ∎
For $n=4$ and $n=5$, large-scale sampling (over 460 million arrangements total) found no counterexamples: every arrangement with $E(S)\approx 0$ (within machine precision) was verified to be a magic square. This computational evidence strongly supports the conjecture that the energy characterization holds for all $n\geq 3$ (see Conjecture A.1 in the Appendix).
@@
For $n=3$ and $n=4$, every magic square is a strict local minimum of $E$ over the discrete space of permutations. That is, for any magic square $S$ and any arrangement $S^{\prime}$ differing from $S$ by a single transposition of two elements, we have $E(S^{\prime})>E(S)=0$.
@@ 3.8 The Moment of Inertia Tensor
The eigenvalues of II—the principal moments of inertia—are invariant under rotation and characterize the intrinsic shape of the mass distribution. By Section 3.6, the off-diagonal elements $I_{xz}=-n^{2}\mathrm{Cov}(X,Z)$ and $I_{yz}=-n^{2}\mathrm{Cov}(Y,Z)$ vanish for magic squares, simplifying the tensor structure.
@@ 4.1 The 3×3 Case: Foundation and Validation
The $3\times 3$ magic square provides the ideal starting point for validating our framework, as it admits a unique solution (up to symmetry) and allows exhaustive analysis. The small size of the arrangement space—9!=362,8809!=362{,}880 permutations—permits complete enumeration, providing definitive verification of the Energy Characterization Theorem.
@@ 4.1.1 Exhaustive Verification of the Energy Characterization
We computed the total magic energy $E(S)$ for all 362,880 permutations of $\{1,\ldots,9\}$. This exhaustive analysis confirms the Energy Characterization Theorem (Section 3.13): an arrangement has zero energy if and only if it is a magic square.
The energy distribution across all permutations (see Section 7) confirms this characterization. The vast majority of arrangements exhibit substantial positive energy, with the distribution spanning from zero to approximately 20. Precisely eight configurations—the Lo Shu and its seven $D_{4}$ symmetry variants—achieve $E(S)=0$ exactly, corresponding to the single essentially different magic square. No false positives (zero energy but non-magic) or false negatives (magic but nonzero energy) exist, providing complete verification of the theorem for $n=3$.
The energy landscape reveals the exceptional nature of magic squares. While generic permutations cluster around a mean energy of approximately 2.5, magic squares are isolated at the absolute minimum. The gap between zero and the smallest positive energy provides a quantitative measure of the “stability” of the magic property (see Section 7 in Section 4.5).
@@ 4.1.2 Perturbation Isolation Analysis
Section 2 presents a comprehensive nine-panel analysis of perturbation gaps across orders $n=3,4,5$, testing all magic squares for each order: all 8 variants of the Lo Shu for $n=3$, all 880 unique $n=4$ magic squares, and 35 representative $n=5$ squares generated via the Siamese method with $D_{4}$ symmetries. Across all 107,388 tested perturbations, every single swap increases the energy, providing strong numerical evidence that magic squares are isolated local minima in the covariance landscape.
@@
Figure 2: Comprehensive Perturbation Analysis Across All Magic Squares. Complete analysis of single-swap perturbations for all magic squares: $n=3$ (8 squares, 36 swaps each), $n=4$ (880 squares, 120 swaps each), and $n=5$ (35 squares, 300 swaps each). Top row: Minimum gap distributions reveal geometric invariance for $n=3$ (all squares identical, $\Delta_{3}=0.0988$ exactly), substantial heterogeneity for $n=4$, and clustering for $n=5$. Middle row: (Left) Top 6 smallest gaps showing high degeneracy—many swaps produce identical energy increases. (Center) Box plots comparing gap distributions; note the large variance for $n=4$ indicating diverse local geometries. (Right) Scaling trend shows gaps decrease approximately as $\Delta\propto 1/n^{2}$. Bottom row: (Left) Cumulative distributions exhibit similar sigmoid forms on log scale. (Center) Gap variation across magic square population reveals three distinct bands for $n=4$, suggesting geometric subclasses with different perturbation resistance. (Right) Summary statistics. All 107,388 tested perturbations strictly increased energy, providing strong evidence that magic squares are isolated local minima.
Section 1 summarizes the perturbation gap statistics. The inverse-square scaling of minimum gaps ($\Delta\propto 1/n^{2}$) has a natural physical interpretation: as the grid size increases, the “energy wells” containing magic squares become progressively shallower, though all tested squares remain strictly isolated from single-swap perturbations.
@@
Table 1: Perturbation gap statistics from comprehensive analysis of all magic squares. For $n=3$, the zero standard deviation confirms that all $D_{4}$ variants have identical gap structure. The minimum gap scales approximately as $\Delta_{\min}\propto 1/n^{2}$.

### v2

@@ Abstract
We introduce Magic Gems, a geometric representation of magic squares as three-dimensional polyhedra. By mapping an $n\times n$ magic square onto a centered coordinate grid with cell values as vertical displacements, we construct a point cloud whose convex hull defines the Magic Gem. Building on prior work connecting magic squares to physical properties such as moment of inertia, this construction reveals an explicit statistical structure: we show that magic squares have vanishing covariances between position and value. We develop a covariance energy functional—the sum of squared covariances with individual row, column, and diagonal indicator variables—and prove that for all $n\geq 3$, an arrangement is a magic square if and only if this complete energy vanishes. This characterization transforms the classical line-sum definition into a statistical orthogonality condition. We also study a simpler “low-mode” relaxation using only four aggregate position indicators; this coincides with the complete characterization for $n=3$ (verified exhaustively) but defines a strictly larger class for $n\geq 4$ (explicit counterexamples computed). Perturbation analysis demonstrates that magic squares are isolated local minima in the energy landscape. The representation is invariant under dihedral symmetry $D_{4}$, yielding canonical geometric objects for equivalence classes.
@@ 1.1 Main Contributions
Our main theoretical contribution is the Complete Energy Characterization Theorem (Theorem 3.17), which establishes that magic squares correspond exactly to zeros of a covariance energy functional using individual row, column, and diagonal indicator variables. This result holds for all $n\geq 3$ and follows from elementary properties of covariances. We also identify a natural four-term “low-mode” relaxation using aggregate position coordinates; this simpler energy characterizes magic squares for $n=3$ but defines a strictly larger class for $n\geq 4$, with explicit counterexamples computed. This result extends the physical interpretation of magic squares initiated by Loly [7], who showed that the moment of inertia tensor of a magic square (treating entries as masses) exhibits special invariance properties. Our covariance formulation makes this connection explicit and enables systematic perturbation analysis, confirming that magic squares are isolated points in the arrangement space.
@@ 1.2 Paper Organization
The remainder of this paper proceeds as follows. Section 2 reviews classical results on magic squares and establishes notation for the geometric and statistical concepts we employ. Section 3 presents the Magic Gem construction, proves the Complete Energy Characterization Theorem, introduces the low-mode relaxation, and develops the perturbation analysis framework. Section 4 describes our computational experiments, validating theoretical predictions and exploring the structure of the arrangement space for $n=3$, 44, and $5$. Section 5 interprets our findings and discusses connections to related areas. Finally, Section 6 summarizes our contributions and outlines directions for future research.
@@ 3.1 From Magic Square to Point Cloud
Figure 1 illustrates the four-step construction process, progressing from the magic square through the centered grid to the final three-dimensional representation, along with all eight $D_{4}$ symmetry variants of the Lo Shu square.
Figure 1: Construction process and $D_{4}$ symmetry structure. Top row (a–d): Construction of a Magic Gem from the Lo Shu square. (a) The Lo Shu magic square—the unique (up to symmetry) $3\times 3$ magic square, with each row, column, and diagonal summing to $M(3)=15$. (b) Placement on a centered coordinate grid with origin at center. (c) Assignment of $z$-values from cell entries. (d) Centered Magic Gem point cloud with convex hull. Bottom section (e–l): The eight $D_{4}$ symmetries of the $3\times 3$ magic square. Each panel shows a compact 2D magic square (top) paired with its corresponding 3D Magic Gem (bottom, large and prominent). The gems are shown from different viewing angles to emphasize their three-dimensional structure. Despite their distinct numerical arrangements, all eight $D_{4}$ variants yield the same 3D Magic Gem polyhedron, demonstrating the geometric invariance of the Magic Gem representation under the symmetry group.
@@ 3.2 The Magic Gem Polyhedron
The construction figure (Figure 1, bottom rows) displays all eight symmetry variants of the $3\times 3$ magic square, illustrating how these numerically distinct arrangements share a common geometric representation.
@@ 3.4.1 Forward Direction: Magic Implies Zero Covariance
3.4.2 On the Insufficiency of Aggregate Covariances
It is natural to ask whether the conditions $\mathrm{Cov}(X,Z)=\mathrm{Cov}(Y,Z)=0$ alone suffice to characterize magic squares, or at least semi-magic squares (those with equal row and column sums but possibly failing the diagonal conditions). Perhaps surprisingly, the answer is no—and understanding why reveals a crucial distinction that motivates our energy functional approach.
Remark 3.8 (Why $\mathrm{Cov}(X,Z)=0$ Is Not Enough).
The condition $\mathrm{Cov}(X,Z)=0$ is equivalent to a single first-moment constraint on column sums. Let $C_{j}=\sum_{i=0}^{n-1}s_{ij}$ denote the sum of column $j$, and let $\delta_{j}=C_{j}-M(n)$ be the deviation from the magic constant. Then $\mathrm{Cov}(X,Z)=0$ is equivalent to
$\sum_{j=0}^{n-1}\left(j-\frac{n-1}{2}\right)\delta_{j}=0.$
Combined with the constraint $\sum_{j}\delta_{j}=0$ (which follows from the total sum being fixed), this leaves an (n2)(n-2)-dimensional space of possible deviation patterns. For $n=3$, the nonzero deviations must have the form $(\delta_{0},\delta_{1},\delta_{2})=t(1,-2,1)$; for $n\geq 4$, the space of “balanced but unequal” column patterns grows rapidly.
Remark 3.9 (Counterexamples Exist for $n=3$).
@@ 3.4.2 On the Insufficiency of Aggregate Covariances
Remark 3.10 (The Four-Term Energy Is Insufficient for $n\geq 4$).
Even more strikingly, adding the diagonal covariance conditions does not resolve the issue for $n\geq 4$. The four covariances $\mathrm{Cov}(X,Z)$, $\mathrm{Cov}(Y,Z)$, $\mathrm{Cov}(D_{\mathrm{main}},Z)$, $\mathrm{Cov}(D_{\mathrm{anti}},Z)$ impose only four linear constraints on the $(n^{2}-1)$-dimensional space of mean-zero value functions. For $n=4$, this leaves an $(n^{2}-5)=11$-dimensional nullspace.
Consequently, there exist $4\times 4$ permutations with all four covariances equal to zero that are not magic squares. For example:
$S=\begin{pmatrix}9&4&12&3\\ 16&1&14&7\\ 10&15&11&8\\ 2&5&6&13\end{pmatrix}$
has both diagonal sums equal to $34$, but row sums $(28,38,44,26)$ and column sums $(37,25,43,31)$. All four covariances vanish, yet the arrangement is not magic.
This analysis reveals that aggregate position coordinates $X$ and $Y$ are too coarse: they capture only the first moment of the row and column sum distributions. A true characterization requires testing against individual row and column indicators, which we develop in Section 3.7.
@@ 3.5 Diagonal Conditions as Covariances
This reformulation expresses the diagonal constraints in the same statistical language as the row/column constraints. Together with the aggregate position covariances $\mathrm{Cov}(X,Z)$ and $\mathrm{Cov}(Y,Z)$, these diagonal covariances define the four-term “low-mode” energy $E_{\mathrm{low}}$ (Definition 3.18). For $n=3$, $E_{\mathrm{low}}(S)=0$ is already sufficient to characterize magic squares (verified by exhaustive enumeration; Proposition 3.19), but for $n\geq 4$ it is only a necessary relaxation (see Remark 3.10). A complete characterization requires testing against individual row and column indicators, developed in Section 3.7.
@@ 3.6 Perturbation Analysis
This result shows that magic squares lie at $\Phi=0$, and that generic perturbations of the $z$-values destroy this balance. Importantly, $\Phi=0$ is not unique to magic squares: already for $n=3$ there exist many non-magic arrangements with $\mathrm{Cov}(X,Z)=\mathrm{Cov}(Y,Z)=0$ (Remark 3.9). Thus $\Phi$ provides a useful first-order diagnostic of balance, but not a complete characterization. The complete energy $E_{\mathrm{full}}$ introduced next resolves this by having zeros exactly at magic squares for all $n\geq 3$ (Theorem 3.17).
@@ 3.7 The Covariance Energy Landscape
As established in Section 3.4.2, the aggregate covariances with $X$, $Y$, and diagonal indicators are insufficient to characterize magic squares for $n\geq 4$. We now develop the complete covariance energy using individual row and column indicators.
3.7.1 Row and Column Indicator Variables
Definition 3.14 (Row and Column Indicators).
For $k,\ell\in\{0,1,\ldots,n-1\}$, define the indicator variables:
$R_{k}(i,j)$ $=\begin{cases}1&\text{if }i=k,\\ 0&\text{otherwise},\end{cases}$ (12)
$C_{\ell}(i,j)$ $=\begin{cases}1&\text{if }j=\ell,\\ 0&\text{otherwise}.\end{cases}$ (13)
These functions select the cells in row $k$ or column $\ell$, respectively.
Proposition 3.15 (Row/Column Covariance Characterization).
For an arrangement $S$ with centered $z$-values:
  1. 1.
    $\mathrm{Cov}(R_{k},Z)=0$ for all $k\in\{0,\ldots,n-1\}$ if and only if all row sums equal $M(n)$.
  2. 2.
    $\mathrm{Cov}(C_{\ell},Z)=0$ for all $\ell\in\{0,\ldots,n-1\}$ if and only if all column sums equal $M(n)$.
@@ 3.7.1 Row and Column Indicator Variables
The mean of $R_{k}$ over all cells is $\bar{R}_{k}=n/n^{2}=1/n$. With $\bar{Z}=0$, we have:
$\mathrm{Cov}(R_{k},Z)=\frac{1}{n^{2}}\sum_{i,j}R_{k}(i,j)\cdot z_{ij}=\frac{1}{n^{2}}\sum_{j=0}^{n-1}z_{kj}.$
Since $z_{kj}=s_{kj}-(n^{2}+1)/2$, the covariance vanishes if and only if $\sum_{j}s_{kj}=n(n^{2}+1)/2=M(n)$, i.e., row $k$ sums to the magic constant. The column case is analogous. ∎
3.7.2 The Complete Magic Energy Functional
Definition 3.16 (Complete Covariance Energy).
For an arrangement $S$, define the complete magic energy as:
$E_{\mathrm{full}}(S)=\sum_{k=0}^{n-2}\mathrm{Cov}(R_{k},Z)^{2}+\sum_{\ell=0}^{n-2}\mathrm{Cov}(C_{\ell},Z)^{2}+\mathrm{Cov}(D_{\mathrm{main}},Z)^{2}+\mathrm{Cov}(D_{\mathrm{anti}},Z)^{2}.$ (14)
This involves 2(n1)+2=2n2(n-1)+2=2n terms (one row/column indicator is redundant due to the fixed total sum).
Theorem 3.17 (Complete Energy Characterization).
For all $n\geq 3$, an arrangement $S$ of $\{1,\ldots,n^{2}\}$ is a magic square if and only if $E_{\mathrm{full}}(S)=0$.
Proof.
Forward direction: If $S$ is a magic square, all rows, columns, and both diagonals sum to $M(n)$. By Proposition 3.15 and Proposition 3.12, all covariances in (14) vanish, hence $E_{\mathrm{full}}(S)=0$.
Converse: If $E_{\mathrm{full}}(S)=0$, then each squared covariance term must be zero (as they are all non-negative). By Proposition 3.15, $\mathrm{Cov}(R_{k},Z)=0$ for $k=0,\ldots,n-2$ implies rows 0 through n2n-2 sum to $M(n)$; since the total sum is fixed at $n^{2}(n^{2}+1)/2=nM(n)$, row n1n-1 also sums to $M(n)$. Similarly, all columns sum to $M(n)$. By Proposition 3.12, both diagonals sum to $M(n)$. Thus $S$ is a magic square. ∎
This theorem establishes a complete characterization: magic squares are precisely the zeros of the energy functional $E_{\mathrm{full}}$, valid for all orders $n\geq 3$. The energy landscape interpretation is now rigorous: magic squares are the unique ground states.
3.7.3 The Low-Mode Relaxation
For comparison, we also define the simpler four-term energy:
Definition 3.18 (Low-Mode Energy).
The low-mode energy uses aggregate position coordinates:
$E_{\mathrm{low}}(S)=\mathrm{Cov}(X,Z)^{2}+\mathrm{Cov}(Y,Z)^{2}+\mathrm{Cov}(D_{\mathrm{main}},Z)^{2}+\mathrm{Cov}(D_{\mathrm{anti}},Z)^{2}.$ (15)
Proposition 3.19 (Low-Mode Characterization).
For $n=3$, $E_{\mathrm{low}}(S)=0$ if and only if $S$ is a magic square (verified by exhaustive enumeration of all 362,880362{,}880 permutations). For $n\geq 4$, the condition $E_{\mathrm{low}}(S)=0$ defines a strictly larger class of “low-mode balanced” arrangements that properly contains the magic squares.
The low-mode energy is computationally simpler and provides a useful first-order diagnostic, but it cannot serve as a characterization for $n\geq 4$. The complete energy $E_{\mathrm{full}}$ is required for the full characterization.
@@ 3.7.3 The Low-Mode Relaxation
For $n=3$ and $n=4$, every magic square is a strict local minimum of $E_{\mathrm{full}}$ over the discrete space of permutations. That is, for any magic square $S$ and any arrangement $S^{\prime}$ differing from $S$ by a single transposition of two elements, we have $E_{\mathrm{full}}(S^{\prime})>E_{\mathrm{full}}(S)=0$.
@@ 3.8 The Moment of Inertia Tensor
The eigenvalues of II—the principal moments of inertia—are invariant under rotation and characterize the intrinsic shape of the mass distribution. By Proposition 3.6, the off-diagonal elements $I_{xz}=-n^{2}\mathrm{Cov}(X,Z)$ and $I_{yz}=-n^{2}\mathrm{Cov}(Y,Z)$ vanish for magic squares, simplifying the tensor structure.
@@ 4.1 The 3×3 Case: Foundation and Validation
The $3\times 3$ magic square provides the ideal starting point for validating our framework, as it admits a unique solution (up to symmetry) and allows exhaustive analysis. The small size of the arrangement space—9!=362,8809!=362{,}880 permutations—permits complete enumeration, providing definitive verification of the low-mode characterization for $n=3$ (Proposition 3.19) and a baseline for the energy landscape viewpoint.
@@ 4.1.1 Exhaustive Verification of the Energy Characterization
We computed the low-mode energy $E_{\mathrm{low}}(S)$ for all 362,880 permutations of $\{1,\ldots,9\}$. This exhaustive analysis confirms the Low-Mode Characterization (Proposition 3.19) for $n=3$: an arrangement has zero low-mode energy if and only if it is a magic square.
The energy distribution across all permutations (see Figure 6) confirms this characterization. The vast majority of arrangements exhibit substantial positive energy, with the distribution spanning from zero to a maximum of about 12 (maximum observed: 12.10). Precisely eight configurations—the Lo Shu and its seven $D_{4}$ symmetry variants—achieve $E_{\mathrm{low}}(S)=0$ exactly, corresponding to the single essentially different magic square. No false positives (zero energy but non-magic) or false negatives (magic but nonzero energy) exist, providing complete verification of the theorem for $n=3$.
The energy landscape reveals the exceptional nature of magic squares. While generic permutations cluster around a mode near 3.12 and have mean energy $\approx 4.44$, magic squares are isolated at the absolute minimum. The gap between zero and the smallest positive energy provides a quantitative measure of the “stability” of the magic property (see Figure 6 in Section 4.5).
@@ 4.1.2 Perturbation Isolation Analysis
Figure 2 presents a comprehensive nine-panel analysis of perturbation gaps across orders $n=3,4,5$, testing all magic squares for $n=3$ and $n=4$ (all 8 variants of the Lo Shu for $n=3$ and all 880 $n=4$ equivalence classes under $D_{4}$), and 35 representative $n=5$ squares generated via the Siamese method with $D_{4}$ symmetries. Across all 116,388 tested perturbations, every single swap increases the energy, providing strong numerical evidence that magic squares are isolated local minima in the covariance landscape.
@@
Figure 2: Perturbation Analysis: Magic Squares as Isolated Minima. Analysis of single-swap perturbations for magic squares using the complete energy $E_{\mathrm{full}}$: $n=3$ (all 8 $D_{4}$ variants, 36 swaps each), $n=4$ (all 880 equivalence-class representatives, 120 swaps each), and $n=5$ (35 representative squares, 300 swaps each). Top row: Minimum gap distributions reveal geometric invariance for $n=3$ (all squares identical, $\Delta_{3}=0.0988$ exactly), substantial heterogeneity for $n=4$, and clustering for $n=5$. Middle row: (Left) Top 6 smallest gaps showing high degeneracy—many swaps produce identical energy increases. (Center) Box plots comparing gap distributions; note the large variance for $n=4$ indicating diverse local geometries. (Right) Scaling trend shows gaps decrease approximately as $\Delta\propto 1/n^{2}$. Bottom row: (Left) Cumulative distributions exhibit similar sigmoid forms on log scale. (Center) Gap variation across magic square population reveals three distinct bands for $n=4$, suggesting geometric subclasses with different perturbation resistance. (Right) Summary statistics. All 116,388 tested perturbations strictly increased energy, supporting the conclusion that magic squares are isolated local minima of $E_{\mathrm{full}}$.
Table 1 summarizes the perturbation gap statistics. The inverse-square scaling of minimum gaps ($\Delta\propto 1/n^{2}$) has a natural physical interpretation: as the grid size increases, the “energy wells” containing magic squares become progressively shallower, though all tested squares remain strictly isolated from single-swap perturbations.
@@
Table 1: Perturbation gap statistics from comprehensive analysis of all $n=3$ and $n=4$ magic squares and a representative sample for $n=5$. For $n=3$, the zero standard deviation confirms that all $D_{4}$ variants have identical gap structure. The minimum gap scales approximately as $\Delta_{\min}\propto 1/n^{2}$.